# Modeling Melt Pool Features and Spatter Using Symbolic Regression and Machine Learning

### v1

@@ Abstract
Additive manufacturing (AM) is a rapidly evolving technology that has attracted applications across a wide range of fields due to its ability to fabricate complex geometries. However, one of the key challenges in AM is achieving consistent print quality. This inconsistency is often attributed to uncontrolled melt pool dynamics, partly caused by spatter which can lead to defects. Therefore, capturing and controlling the evolution of the melt pool is crucial for enhancing process stability and part quality. In this study, we developed a framework to support decision-making in AM operations, facilitating quality control and minimizing defects via machine learning (ML) and polynomial symbolic regression models. We implemented experimentally validated computational tools as a cost-effective approach to collect large datasets from laser powder bed fusion (LPBF) processes. For a dataset consisting of 281 process conditions, parameters such as melt pool dimensions (length, width, depth), melt pool geometry (area, volume), and volume indicated as spatter were extracted. Using machine learning (ML) and polynomial symbolic regression models, a high R2 of over 95 % was achieved in predicting the melt pool dimensions and geometry features for both the training and testing datasets, with either process conditions (power and velocity) or melt pool dimensions as the model inputs. In the case of volume indicated as spatter, R2 improved after logarithmic transforming the model inputs, which was either the process conditions or the melt pool dimensions. Among the investigated ML models, the ExtraTree model achieved the highest R2 values of 96.7 % and 87.5 %. With respect to the symbolic regression model, R2 values of 85 % and 82 % were achieved for the training and testing datasets, respectively. Our study culminated in the discovery of symbolic equations based on model inputs in the polynomial fitting model for all investigated parameters, thereby providing interpretable evaluations of the feature importance.
@@ 1 Introduction
Laser Powder Bed Fusion (LPBF) is a popular metal additive manufacturing (AM) process that has demonstrated high precision and remarkable performance in producing complex part geometries and internal structures with micron-scale precision. The operating process of Laser Powder Bed Fusion (LPBF) involves guiding a laser using a scanning galvanometer to selectively melt a layer of powder deposited on a build plate. Once solidification occurs, the next layer of powder is spread and the process is repeated iteratively until the final part is complete [1]. In addition, LPBF offers significant advantages, including design flexibility, an automated method for building parts directly at their location of use from CAD files, and reduced material waste. These benefits have contributed to its adoption in various industries, including aerospace, automotive, medicine, energy, and prototyping [2, 3]. However, despite its increasing application in different sectors, compared to other forms of manufacturing processes such as rolling, casting, and machining, LPBF still occupies a low market share, which is attributed to its low rate of part production and poor quality control. A popular approach to remediate low part production involves using multi-lasers, which, however, increases the defect formation rate. The efficacy of state-of-the-art methods for quality control is determined by comparing the properties and performance of the finished part with the target standard, whereby variability and suboptimal performance are still reported.
@@
Identifying optimal process conditions typically relies on costly and time-intensive trial-and-error experiments across a wide range of process parameters. This challenge has motivated the use of 3D numerical modeling tools, which can effectively capture the complex phenomena occurring during melting and provide insights into parameters that are difficult or impossible to measure using current state-of-the-art in situ experimental systems. However, a high-fidelity AM simulation package that will account for all of the complex physics in AM is computationally expensive. As a result, increasing attention is being directed towards a robust smart AM manufacturing tool, which has been dubbed a solution to match the desired properties and performance of parts and AM tools [4]. This approach entails developing a digital twin of the LPBF processes which will possess the capability to mirror the real-time evolution of the part-building processes with the display of the key measurements and also for-cast future events during laser-material interactions. This type of system will enhance the detection of faults or anomalies during the printing process and enable the development of a sensitive feed-forward or backward control system to mitigate or prevent the intricacies that may occur during the printing process.At the core of the digital twin will be an integration of experimental results, system modeling analysis, and a ML model. ML enables predictive and performance insights that can be difficult to uncover through traditional physics-based modeling approaches due to the high computation cost. [5, 6, 7, 8]. Monitoring the formation and dynamic processes that occur in the melt pool is one of the crucial aspects of LBPF that can improve quality control by providing an understanding of the underlying processes taking place in relation not only to the process conditions, but also to the material types that govern the nature of the melt pool, microstructure, and mechanical properties [2]. Consequently, defect formation, such as porosity, surface roughness, residual stresses, warping, cracking, and delamination that can influence quality control, will be further elucidated and prevented.
@@
The general characteristic of the melt pool that has been popularly monitored to understand the formation of defects or the properties of built parts is dimensions [9, 10, 11, 12, 13], shape[14, 15, 16, 17, 18], temperature distribution [19, 20, 21, 22, 23, 24, 25], oxygen content [26, 27, 28, 29, 30], and pressure [28]. Many studies have been reported on coupling ML with monitoring parameters collected either through experimental or modeling methods as a reduced-order method. Xiao et al. [4] predicted the future area of the melt pool using the collection of parts built from the past history by initially filtering out the noise through a a convolutional neural network (CNN)-based model, then quantitatively predicted the area of future melt pools, using artificial neural network (ANN) trained with scanning history, particularly including past melt pools as inputs from built parts, which significantly reduced the average relative error magnitude (AREM) to 2.8 %, compared to the 14.8 % AREM of the existing Neighboring Effect Modeling Method (NBEM). Akbari et al. [31] predicted the dimensions of the melting pool using ML models, using process parameters and material properties curated from different source of published literatures. Zhang et al. [32] implemented LSTM-based approach to estimate the area of the meltpool with accuracy of 90.7 %, then used Melt Pool Generative Adversarial Network synthesizing the images of the melt pool and achieve a structural similarity score of 0.91. Wang et al. [33] implemented a machine learning-assisted approach based on a deep neural network to demodulate the optical signal to thermal distribution and significantly improve spatial resolution to 28.8 µm/pixel spatial resolution and 10 kHz sampling frequency, ideal for measuring the sharp thermal gradient and cooling rates in the L-PBF process. Scime et al. [34] used a high-speed camera to capture the molten pool images. An unsupervised learning algorithm was proposed to distinguish the molten pool and identify defects.
@@
Extending quality control to the monitoring of the spatter ejection from the melt pool can provide information on the state or stability of the melt pool as defects develop [35, 36, 37, 38]. In-homogeneity fusion and thermal gradient caused by spatter can be a source of crack initiation sites [39]. The traveling spatter can absorb the laser radiation or reflect it, causing a fluctuation in the intensity of the laser [40]. Limited studies have been conducted on augmenting ML models and spatter formation, and on this path, most of the work has focused on spatter detection by coupling ML with computer vision or classifying according to process conditions or signals [3, 41, 42, 43, 44, 45], which does not provide the underlying mechanisms to the ejection rate of spatter as a means of quality control. Previously, we implemented ML to classify and detect the slight variation in the measured parameters between the spatter and the melt pool at 5 µs, unveiling the ranking of the importance of different measured parameters towards spatter ejection. Complemented by an explainable AI technique known as SHapley Additive exPlanations (SHAP), we identify the range of feature values that positively contribute to spatter predictions [5].
@@
Expanding the boundaries of our previous investigations, in this work we conduct both machine learning (ML) regression and symbolic regression to predict melt pool features measured at a stable state, such as dimensions (length, width, and depth) and geometric properties (cross-sectional area and volume), as well as the volume indicated by spatter. We collected our data set of spatter count versus process conditions of different power/velocity from augmentation of OpenFOAM and FLOW-3D simulations via ML classification tasks as a reduced-order method for spatter predictions. The motivation behind the techniques lies on the trade-off between the capability and computational costs of both simulation tools. However, FLOW-3D incorporates certain restrictive assumptions to reduce computational costs, which limit its ability to simulate spatter accurately. Specifically, it approximates the keyhole gas dynamics using estimates of recoil pressure and mass transfer, rather than directly simulating them. Furthermore, FLOW-3D does not account for the tangential surface tension component [46, 47]. In contrast, our OpenFOAM model directly simulates keyhole gas dynamics and incorporates tangential surface tension. The analysis indicated that the melt pool features can be predicted with a high R2 value from the process conditions (power and velocity). Similarly, the volume indicated as spatter can be predicted with a high R2 value from either the process conditions or the melt pool dimensions, as shown in Fig. 1. Our findings not only demonstrate the interconnection between process conditions, melt pool features, and spatter counts but also provide an interpretable mathematical expression of the parameters derived from the input variables through symbolic regression models, which could be crucial for process optimization.

### v2

# Discovery of Spatter Constitutive Models in Additive Manufacturing Using Machine Learning

@@ Abstract
Additive manufacturing (AM) is a rapidly evolving technology that has attracted applications across a wide range of fields due to its ability to fabricate complex geometries. However, one of the key challenges in AM is achieving consistent print quality. This inconsistency is often attributed to uncontrolled melt pool dynamics, partly caused by spatter which can lead to defects. Therefore, capturing and controlling the evolution of the melt pool is crucial for enhancing process stability and part quality. In this study, we developed a framework to support decision-making towards efficient AM process operations, capable of facilitating quality control and minimizing defects via machine learning (ML) and polynomial symbolic regression models. We implemented experimentally validated computational tools, specifically for laser powder bed fusion (LPBF) processes as a cost-effective approach to collect large datasets. For a dataset consisting of 281 varying process conditions, parameters such as melt pool dimensions (length, width, depth), melt pool geometry (area, volume), and volume indicated as spatter were extracted. Using machine learning (ML) and polynomial symbolic regression models, a high R2 of over 95 % was achieved in predicting the melt pool dimensions and geometry features on both the training and testing datasets, with either process conditions (power and velocity) or melt pool dimensions as the model inputs. In the case of volume indicated as spatter, the value of the R2 improved after logarithmic transforming the model inputs, which were either the process conditions or the melt pool dimensions. Among the investigated ML models, the ExtraTree model achieved the highest R2 values of 96.7 % and 87.5 %. With respect to the symbolic regression model, R2 values of 85 % and 82 % were achieved on the training and testing datasets, respectively. Our study culminated in the discovery of symbolic equations based on model inputs in the polynomial fitting model for all investigated parameters, thereby providing interpretable evaluations of the feature importance.
@@ 1 Introduction
Laser Powder Bed Fusion (LPBF) is a popular metal additive manufacturing (AM) process that has demonstrated high precision and remarkable performance in producing complex part geometries and internal structures with micron-scale precision. The operating process of Laser Powder Bed Fusion (LPBF) involves guiding a laser using a scanning galvanometer to selectively melt a layer of powder deposited on a build plate. Once solidification occurs, the next layer of powder is spread and the process is repeated iteratively until the final part is complete [1]. In addition, LPBF offers significant advantages, including design flexibility, an automated method for building parts directly at their point of use from CAD files, and reduced material waste. These benefits have facilitated its adoption in various industries, including aerospace, automotive, medicine, energy, and prototyping [2, 3]. However, despite its increasing application in different sectors, compared to other forms of manufacturing processes such as rolling, casting, and machining, LPBF still occupies a low market share, which is attributed to its low rate of part production and poor quality control. A popular approach to remediate low part production involves using multi-lasers, which, however, increases the defect formation rate. The efficacy of state-of-the-art methods for quality control is assessed by comparing the properties and performance of the finished part with the target standard, although variability and suboptimal performance are still observed.
@@
Identifying optimal process conditions typically relies on costly and time-intensive trial-and-error experiments across a wide range of process parameters. This challenge has motivated the use of 3D numerical modeling tools, which can effectively capture the complex phenomena occurring during melting and provide insights into parameters that are difficult or impossible to measure using current in situ experimental techniques. However, a high-fidelity AM simulation package that will account for all of the complex physics in AM is computationally expensive. As a result, increasing attention is being directed towards a robust smart AM manufacturing tool, which has been dubbed a solution to match the desired properties and performance of parts and AM tools [4]. This approach involves developing a digital twin of the LPBF process, which will be capable of mirroring the real-time evolution of part-building processes, displaying key measurements, and forecasting future events during laser-material interactions. This type of system will enhance the detection of faults or anomalies during the printing process through the integration of a sensitive control system to mitigate or prevent the intricacies that may occur during the printing process. The core component of the digital twin will be the integration of experimental datasets, system modeling analysis, and ML models. ML enables predictive and performance insights that can be difficult to uncover through traditional physics-based modeling approaches due to the high computation cost. [5, 6, 7, 8, 9]. Monitoring the formation and dynamic processes that occur in the melt pool is one of the crucial aspects of LBPF operation that can improve quality control by providing an understanding of the underlying processes that take place in relation not only to the process conditions, but also to the material types that govern the nature of the melt pool, the microstructure, and the mechanical properties [2]. Consequently, defect formation, such as porosity, surface roughness, residual stresses, warping, cracking, and delamination that can influence quality control, will be further elucidated and prevented.
@@
The general characteristics of the melt pool that have been popularly monitored to understand the formation of defects or the properties of built parts are dimensions [10, 11, 12, 13, 14], shape[15, 16, 17, 18, 19], temperature distribution [20, 21, 22, 23, 24, 25, 26], oxygen content [27, 28, 29, 30, 31], and pressure [29]. Many studies have been reported on coupling ML with monitoring parameters collected either through experimental or modeling methods to accelerate the interpretation of system behavior. Xiao et al. [4] predicted the future melt pool area using the history of the previous construction. They first filtered out noise using a convolutional neural network (CNN)-based model, then quantitatively estimated future melt pool areas using an artificial neural network (ANN) trained on scanning history, particularly past melt pools. This approach significantly reduced the average relative error magnitude (AREM) to 2.8% compared to 14.8% with the existing Neighboring Effect Modeling Method (NBEM). Akbari et al. [32] predicted the dimensions of the melting pool using ML models, where the process parameters and the material properties are curated from different source of published literatures. Zhang et al. [33] implemented LSTM-based approach to estimate the area of the meltpool with accuracy of 90.7 %, then used Melt Pool Generative Adversarial Network to synthesize the images of the melt pool and achieve a structural similarity score of 0.91. Wang et al. [34] implemented a machine learning-assisted approach based on a deep neural network to demodulate the optical signal to thermal distribution and significantly improve spatial resolution to 28.8 µm/pixel spatial resolution and 10 kHz sampling frequency, ideal for measuring the sharp thermal gradient and cooling rates in the L-PBF process. Scime et al. [35] used a high-speed camera to capture the images of the molten pool, then proposed an unsupervised learning algorithm to distinguish the molten pool and identify defects.
@@
Extending quality control to the monitoring of the ejection of the spatter from the melt pool can provide information on the state or stability of the melt pool as defects develop [36, 37, 38, 39]. In-homogeneity fusion and thermal gradient caused by spatter can be a source of crack initiation sites [40]. The traveling spatter can absorb the laser radiation or reflect it, causing a fluctuation in the intensity of the laser [41]. Limited research has explored the integration of machine learning (ML) models with spatter formation. Most studies in this area have focused mainly on spatter detection by coupling ML with computer vision or by classifying spatter according to process conditions or signals [3, 42, 43, 44, 45, 46]. However, these approaches do not explain the underlying mechanisms that govern the ejection rate of the spatter, which fluctuates according to the process conditions and the state of the melt pool. Previously, we implemented machine learning (ML) to classify and detect slight variations in the measured parameters between the spatter and the melt pool at 5 µs. This approach revealed the ranking of different measured parameters based on their importance in spatter ejection. To enhance interpretability, we applied an explainable AI technique known as SHapley Additive exPlanations (SHAP), which allowed us to identify the range of the feature values that positively contribute to spatter predictions [5].
@@
Building upon our previous investigations, this work employs both ML regression and symbolic regression to predict melt pool features measured at a stable state. These features include dimensions (length, width, and depth), geometric properties (cross-sectional area and volume), and the volume indicated by spatter. We collect the data set of the spatter count versus process conditions by augmenting OpenFOAM and FLOW-3D simulations via ML model developed for binary classification task for spatter/melt pool predictions. The motivation behind the techniques lies on the trade-off between the capability and computational costs of both simulation tools. FLOW-3D incorporates certain restrictive assumptions to reduce computational costs, which limit its ability to simulate spatter accurately. Specifically, it approximates the keyhole gas dynamics using estimates of recoil pressure and mass transfer, rather than directly simulating them. Furthermore, FLOW-3D does not account for the tangential surface tension component [47, 48]. In contrast, our OpenFOAM model directly simulates keyhole gas dynamics and incorporates tangential surface tension. The analysis indicated that the melt pool features can be predicted with a high R2 value from the process conditions (power and velocity). Similarly, the volume indicated as spatter can be predicted with a high R2 value from either the process conditions or the dimensions of the melt pool, as shown in Fig. 1. Our findings not only demonstrate the interconnection between process conditions, melt pool features, and spatter counts, but also provide an interpretable mathematical expression of the parameters derived from the input variables through symbolic regression models, which could be crucial for process optimization.